The vapour pressure of a pure substance at a temparature is 40 bar. the actuel and ideal gas values of g/RT for tha saturated vapour at this temparature T and 40 bar are 5.0 and 5.8 respectevly. The figacity of the saturated liquid at these conditions is
1
Expert's answer
2020-07-29T06:32:38-0400

Given : P = 40 bar "G\/RT" = 5.0 Gig/RT  =5.8

We kow that   GR RTln(fi/P) ⟹ (G − Gig) =RTln(fi/P) ⟹

(G/RT - GR/RT) =ln(fi/P) ⟹ (5.0 −5.8)= ln(fi/P)  ⟹

ln(fi/P) = −0.8 ⟹ fi/P = exp−0.8  ⟹ fi = P x exp−0.8  = 40×exp−0.8 = 17.98 bar